OpenStudy (anonymous):

What is the radius of a circle that has a circumference of 81.64 meters?

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

This is the equation: 81.64=2(3.14)r

OpenStudy (anonymous):

so 2 would be the answer ?

OpenStudy (pinksapphire):

No, the radius is going to be 13.

OpenStudy (pinksapphire):

Do you want me to show you how to do it?

OpenStudy (anonymous):

how u get 13

OpenStudy (anonymous):

yes plz

OpenStudy (pinksapphire):

81.64=2(3.14)r Remember that r is the radius. 2*3.14 Multiply 2 by 3.14 (3.14 is pi). 81.64=6.28r So now you have this... 81.64/6.28=6.28/6.28r Now you divide 6.28 to get r by itself. 13=r or r=13

OpenStudy (anonymous):

still don't get it lol

OpenStudy (pinksapphire):

Um...did you learn this yet?

OpenStudy (anonymous):

yes but i don't understand it

OpenStudy (anonymous):

What is the diameter of a circle that has a circumference of 56.52 feet? 56.52=2(3.14)r 3.14 times 2=6.28 56.52=6.28 ----- --- 6.28 6.28 ---- ---- 9 1

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

so r=9

OpenStudy (pinksapphire):

No, that would be the radius...but the diameter would be 18.

OpenStudy (anonymous):

how would u find the diameter

OpenStudy (pinksapphire):

Radius is half of the diameter.

OpenStudy (anonymous):

how u get 18 do u have to divide 9 by 3.14?

OpenStudy (pinksapphire):

No, you multiply the radius by 2. 9*2=18

OpenStudy (anonymous):

What is the circumference of a circle that has a radius of 10 centimeters? do u do the same thing for the circumference ?

OpenStudy (anonymous):

so 10=2(3.14)r 10=6.28 10=6.28 --- ----- 6.28 6.28 ---- ---- 1.59 1 ---- ----- 1.59 r=1.59 1.59times 2=3.18

OpenStudy (pinksapphire):

To find the circumference, you have to multiply the diameter and 3.14.

OpenStudy (anonymous):

the diameter is gonna be 10 right

OpenStudy (anonymous):

so 10times 3.14 is 31.4

OpenStudy (pinksapphire):

10 is the radius. 20 is the diameter. 20(3.14)=62.8

OpenStudy (anonymous):

how is 20 the diameter

OpenStudy (pinksapphire):

10 is half the distance of the circle. 20 is the full distance across the circle.